the measure of one of two supplementary angles is 4 greater than 3 times the other. Find the measure of the larger angle
Answer by mananth(16946)

let angle be x
supplement will be 180-x
x=3(180-x)+4
x= 540-3x+4
x+3x=544
4x=544
x=136 deg the larger angle